# LLM-Oriented Information Retrieval: A Denoising-First Perspective

### 01 May 2026 v1

@@ 3.4. Retrieval Verification
(1) Evaluating Utility of Retrieval. Effective denoising requires evaluating whether retrieved context genuinely benefits the generation, rather than merely relevance matching in an LLM-oriented system. Traditional metrics like NDCG fail to capture the nuanced relationship between retrieval and generation quality (Cuconasu et al., 2024). To disentangle these effects, LLM-as-a-judge frameworks such as RAGAS (ES et al., 2023) and RAGChecker (Ru et al., 2024) separately measure components such as retrieval quality, context precision and answer faithfulness, making it possible to identify system failure modes. Furthermore, to isolate the true information gain, utility-based metrics like SePer (Dai et al., ) quantify the exact reduction in semantic uncertainty provided by the retrieval stage to measure retrieval utility. Complementary benchmarks such as RAGTruth (Niu et al., 2024), FaithEval (Ming et al., ), SafeRAG (Liang et al., 2025b), and RGB (Chen et al., 2024b) stress whether systems preserve faithfulness when evidence density degrades.
@@ 3.5. Closed-Loop Training
Agentic loop orchestration and regulation. To mitigate cumulative noise, agents employ structured flow control as active noise filters. First, query decomposition minimizes semantic drift: methods like IRCoT (Trivedi et al., 2023) and ChainRAG (Zhu et al., 2025a) break complex queries into verifiable sub-steps to prevent “lost-in-retrieval” failures, while hierarchical planners in LevelRAG (Zhang et al., 2025f) and MindSearch (Chen et al., ) ensure narrowly targeted retrieval to minimize distractor intrusion. Second, adaptive scheduling acts as a dynamic gate: systems like Self-RAG (Asai et al., ), FLARE (Jiang et al., 2023b), and DRAGIN (Su et al., 2024) assess parametric uncertainty to decide when to retrieve, blocking unnecessary contexts that would dilute focus. Finally, orchestration must incorporate adversarial filtering against prompt injection (Liu et al., 2024c) and corpus poisoning (Zou et al., 2025; Shafran et al., 2025), requiring interface-level sanitization such as structured queries (Chen et al., 2025).
@@
S4.3 Deep research and reports R1 subquery drift; A1 evidence dilution; V1 attribution gap. Symptom: confident claims with weak/irrelevant citations. Query: “Compare PFAS regulations US vs EU (2025).” Subqueries drift to general health-effects; report claims “banned globally” with a citation to an unrelated blog snippet. RR subquery anchoring + novelty/duplicate filter. CA claim-level evidence packing (“cite-by-claim”). FV entailment check between each claim and cited spans. CL RL-trained search policies optimize evidence density. JUDGE RACE (Overall, Comprehensiveness, Depth, Instruction-Following, Readability). CITE FACT (Citation Accuracy,Effective Citation) (deepresearchbench); UTILITY SePer (Dai et al., ). CITE Citation Precision/Recall (Gao et al., 2023b).
@@ 4.1. Coding Agents
Finally, coding agents uniquely employs execution and tests as a verification tool. The edit-execute loop allows agents to dynamically prune noisy trajectories using precise feedback from failing tests and stack traces. This signal enables systems like SWE-agent to distinguish plausible hallucinations from functional solutions, iteratively refining the retrieval context (Yang et al., 2024; Antoniades et al., 2025). Moreover, recent works suggest that strengthening this oracle via test augmentation or differential testing further prevents overfitting to sparse supervision (Yu et al., 2025; Wang et al., 2025c). Consequently, retrieval pipeline in coding agents becomes a cyclic, verifiable loop where code structure and execution states continuously gate the flow of information.

### 17 May 2026 v2

@@ 3.4. Retrieval Verification
(1) Evaluating Utility of Retrieval. Effective denoising requires evaluating whether retrieved context genuinely benefits the generation, rather than merely relevance matching in an LLM-oriented system. Traditional metrics like NDCG fail to capture the nuanced relationship between retrieval and generation quality (Cuconasu et al., 2024). To disentangle these effects, LLM-as-a-judge frameworks such as RAGAS (ES et al., 2023) and RAGChecker (Ru et al., 2024) separately measure components such as retrieval quality, context precision and answer faithfulness, making it possible to identify system failure modes. Furthermore, to isolate the true information gain, utility-based metrics like SePer (Dai et al., 2025) quantify the exact reduction in semantic uncertainty provided by the retrieval stage to measure retrieval utility. Complementary benchmarks such as RAGTruth (Niu et al., 2024), FaithEval (Ming et al., ), SafeRAG (Liang et al., 2025b), and RGB (Chen et al., 2024b) stress whether systems preserve faithfulness when evidence density degrades.
@@ 3.5. Closed-Loop Training
Agentic loop orchestration and regulation. To mitigate cumulative noise, agents employ structured flow control as active noise filters. First, query decomposition minimizes semantic drift: methods like IRCoT (Trivedi et al., 2023) and ChainRAG (Zhu et al., 2025a) break complex queries into verifiable sub-steps to prevent “lost-in-retrieval” failures, while hierarchical planners in LevelRAG (Zhang et al., 2025f) and MindSearch (Chen et al., ) ensure narrowly targeted retrieval to minimize distractor intrusion. Second, adaptive scheduling acts as a dynamic gate: systems like Self-RAG (Asai et al., ), FLARE (Jiang et al., 2023b), and DRAGIN (Su et al., 2024a) assess parametric uncertainty to decide when to retrieve, blocking unnecessary contexts that would dilute focus. Finally, orchestration must incorporate adversarial filtering against prompt injection (Liu et al., 2024c) and corpus poisoning (Zou et al., 2025; Shafran et al., 2025), requiring interface-level sanitization such as structured queries (Chen et al., 2025).
@@
S4.3 Deep research and reports R1 subquery drift; A1 evidence dilution; V1 attribution gap. Symptom: confident claims with weak/irrelevant citations. Query: “Compare PFAS regulations US vs EU (2025).” Subqueries drift to general health-effects; report claims “banned globally” with a citation to an unrelated blog snippet. RR subquery anchoring + novelty/duplicate filter. CA claim-level evidence packing (“cite-by-claim”). FV entailment check between each claim and cited spans. CL RL-trained search policies optimize evidence density. JUDGE RACE (Overall, Comprehensiveness, Depth, Instruction-Following, Readability). CITE FACT (Citation Accuracy,Effective Citation) (deepresearchbench); UTILITY SePer (Dai et al., 2025). CITE Citation Precision/Recall (Gao et al., 2023b).
@@ 4.1. Coding Agents
Finally, coding agents uniquely employs execution and tests as a verification tool. The edit-execute loop allows agents to dynamically prune noisy trajectories using precise feedback from failing tests and stack traces. This signal enables systems like SWE-agent to distinguish plausible hallucinations from functional solutions, iteratively refining the retrieval context (Yang et al., 2024; Antoniades et al., 2025). Moreover, recent works suggest that strengthening this oracle via test augmentation or differential testing further prevents overfitting to sparse supervision (Yu et al., 2025a; Wang et al., 2025d). Consequently, retrieval pipeline in coding agents becomes a cyclic, verifiable loop where code structure and execution states continuously gate the flow of information.